OpenStudy (anonymous):

What is the x-intercept of the line with the equation 1/3x + y = -15 ? ( ? , 0)

OpenStudy (secret-ninja):

Try to get y by itself on one side of the equal sign. Do you know how?

OpenStudy (anonymous):

do i cancel out 1

OpenStudy (secret-ninja):

One way to do it to just get the 1/3x and put it on the right side. You just have to change the symbol, and since it has none, you can assume it is positive, so you change it to negative. Your result: \[y = -\frac{ 1 }{ 3 } x -15\]

OpenStudy (secret-ninja):

The equation for a line on graph is: y=mx+b The m is the slope (in this case, the -1/3x) and the b is the y intercept.

OpenStudy (secret-ninja):

Knowing this, what do you think the y intercept is?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on

OpenStudy (secret-ninja):

Sure. :)

OpenStudy (secret-ninja):

Would you like for me to show you?

OpenStudy (anonymous):

lol yes please

OpenStudy (secret-ninja):

Okay. ^-^ As I said before: "The equation for a line on graph is: y=mx+b The m is the slope (in this case, the -1/3x) and the b is the y intercept." Catch that? The "b is the y-intercept"...

OpenStudy (secret-ninja):

What number is in the b's place?

OpenStudy (anonymous):

-15

OpenStudy (secret-ninja):



OpenStudy (secret-ninja):

Correct! Good job. :D

OpenStudy (asnaseer):

@Secret-Ninja - the question is asking for the "x-intercept" not the "y-intercept"

OpenStudy (secret-ninja):

:O

OpenStudy (secret-ninja):

OMG

OpenStudy (secret-ninja):

*FACEPALMS REALLY HARD*

OpenStudy (asnaseer):

@CJBALLER5323 - when a curve or line touches the x-axes what is the value of y?

OpenStudy (secret-ninja):

I'm so stupid.

OpenStudy (anonymous):

lol its ok

OpenStudy (asnaseer):

@Secret-Ninja - no you are not - we all make mistakes - that is what makes us human :)

OpenStudy (anonymous):

@asnaseer im not sure

OpenStudy (secret-ninja):

Curse my terrible eyesight.. At least now you know how to find the y intercept.. hehe...

OpenStudy (anonymous):

3/4?

OpenStudy (asnaseer):

let me draw a graph to help you @CJBALLER5323 ...

OpenStudy (asnaseer):